Alex has twice as much money as Jennifer. Jennifer has $6 less than Shannon . All together the three friends have $54. How much money does each have?
Answer by ankor@dixie-net.com(22740)

Alex has twice as much money as Jennifer.
a = 2j
:
Jennifer has $6 less than Shannon.
j = s-6
j+6 = s
;
All together the three friends have $54.
a + j + s = 54
Replace a and s
2j + j + (j+6) = 54
4j = 54 - 6
4j = 48
j = 48/4
j = 12 is Jen's age
:
You can find the ages of the other two, ensure they add up to 54
: